OpenStudy (anonymous):

@speedygonzalles can you help me please Let f(x) = x^2 - 81 find f^1(x)

jigglypuff314 (jigglypuff314):

to find the inverse of a function replace x with y and y with x so x = y^2 - 81 solve for y

OpenStudy (anonymous):

would i add 81 to y^2?

OpenStudy (anonymous):

?

jigglypuff314 (jigglypuff314):

mmm if by add you mean x = y^2 - 81 +81 +81 then yes :P

OpenStudy (anonymous):

so it would be 81x =y^2 do i divide 81 by 2 now?

jigglypuff314 (jigglypuff314):

not quite :) what I did was addition so you would get x + 81 = y^2 so you would then square root both sides

OpenStudy (anonymous):

so it would be x + 9 = y^2?

OpenStudy (anonymous):

i dont know what to do with y^2

jigglypuff314 (jigglypuff314):

square rooting both sides would make y^2 turn into y and (x+81) with a square root over it

OpenStudy (anonymous):

would that make it ±sqrt x plus 81?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

thank you so much can you help me with a couple more?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (anonymous):

Thank you! my next problem is Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (anonymous):

i think i would do f(g(2)) = x + 8 is that right?

jigglypuff314 (jigglypuff314):

not quite :) first you should find g(2) = ?

OpenStudy (anonymous):

ok how would i do that is it f(g(2)) = x + 8 g(x) = 2x - 6x -7?

OpenStudy (anonymous):

im not sure how to set the problem up im sorry

jigglypuff314 (jigglypuff314):

is this the same problem? :/

OpenStudy (anonymous):

yes its my 2nd question its not the same as the sqrt81 problem

jigglypuff314 (jigglypuff314):

I was getting a little confused because "f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2))." was not the same as "f(g(2)) = x + 8 g(x) = 2x - 6x -7?"

OpenStudy (anonymous):

oh im sorry i didnt mean to confuse you

jigglypuff314 (jigglypuff314):

mmm alright the over all method to doing a f(g(n)) problem would be to first plug in n (number) into g(x) to get another number which you would then plug in to f(x) and simplify to get your answer

OpenStudy (anonymous):

but i dont know what to plug into my problem

jigglypuff314 (jigglypuff314):

well they want you to find f(g(2)) right? so then plug in 2 into g(x)

OpenStudy (anonymous):

would it be f(x + 8) = g(x2 - 6 -7)

jigglypuff314 (jigglypuff314):

whoa there... is it "f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2))." that you're looking for? then you just need to find g(2) first

OpenStudy (anonymous):

yes that is my problem so since i only have to find g(2) would i do g(x2 -6 - 7)?

jigglypuff314 (jigglypuff314):

not quite :) g(2) means to plug in 2 for x into x^2 - 6x - 7

OpenStudy (anonymous):

ok:) so would it be x^2 -6x -7? and if so would my next step be +7 to both sides? im sorry i hope im not being to much of a pain algebra for some reason is really hard for me to grasp

jigglypuff314 (jigglypuff314):

it's alright :) mmm do you know what to plug in 2 for x means? :/

OpenStudy (anonymous):

does it mean it is now 2x?

OpenStudy (anonymous):

or would it mean x = 2

jigglypuff314 (jigglypuff314):

yes x = 2 is what I mean :)

OpenStudy (anonymous):

so would my equation be 2x^2 -6x -7?

jigglypuff314 (jigglypuff314):

mmm more like (2)^2 - 6(2) - 7 and simplify using PEMDAS

OpenStudy (anonymous):

ok:) can you tell me if this is right? (2)^2 - 6(2) - 7 4 - 12 -7 +4 +4 -12 -3 / 3 = -4?

jigglypuff314 (jigglypuff314):

not quite :) you're over thinking this :P just do 4 - 12 - 7 like a calculator would (sorry but my internet died, so I might be slower at responding)

OpenStudy (anonymous):

-15?

jigglypuff314 (jigglypuff314):

Correct! :) then plug in -15 into x+8 so -15 + 8 = ?

OpenStudy (anonymous):

-7?

jigglypuff314 (jigglypuff314):

Yes! :D tah-dah! ^_^

OpenStudy (anonymous):

Thank you so much!!! can i open a new question and tag you in it?

jigglypuff314 (jigglypuff314):

sure :)